Firewall upgrade is complete and no issues encountered.
We will be upgrading our firewall early Friday morning to incorporate a new security patch. The actual down time should be mininal, because a reboot is required to implement the new software.
Updated 9/25 11:51 AM
We currently are experiencing problems with www2.ups.edu. FTP service, which was not working earlier, has been restored.
We plan to reboot this system tomorrow morning (9/26/08) before 9:00 AM and hope that this will resolve the issues we’re having. More will posted here as we continue investigate the problems.
On Sunday the 21st of September 2008, there was scheduled maintenance on Camano and Grace, for Solaris9 cluster patches as well as J2SE “java” related patches. There was also a hardware replacement of a power supply for the server known as Camano. Both servers were up and running without incident by 10:30 a.m. Sunday.
Library.ups.edu has been experiencing slow response times and dropped connections for several days. We have made some changes to the way the server responds and now are waiting to see if this resolves the problem.
Our ISP, had a router issue which caused our internet access to be down during the night. The service was restored at 6:57am this morning.
Service has been restored to all central database systems (Cascade, Cascade Web, Banner, Famis, Millennium, Basis and CRM) as of 3pm. Please contact the Help Desk, 253.879.8585 or firstname.lastname@example.org, if you continue to experience problems.
All central database systems are unavailable. This service disruption began at 11AM today and services impacted include: Cascade, Cascade Web, Banner, Famis, Basis, CRM and Millennium. Technology Services has determined the source of the disruption to be a power outage and is working with Facilities Services to resolve as quickly as possible.
UPDATE AS OF 2PM:
Stable power has been restored and Technology Services is now focused on restarting and verifying the database systems. Every effort is being made to restore services as quickly as possible. Our earliest continuation of service is estimated at 3pm today.
At 6 AM this morning, a reload of the core switch caused a failure of the network interfaces on the Cascade Web server. This in turn caused the Cascade web interface to be unavailable from 6 to 8:15 this morning. The server was rebooted and recovered, restoring service at approximately 8:15.
Other affected servers were grace, camano, and crystal, all of which were rebooted to restore service at 7:30 AM. Batch job processing was affected between 6 and 7:30 AM.